Rational Demiconductor has the following market-value balance sheet:

Rational Demiconductor Balance Sheet (Market Values, in $ millions)

Surplus cash

$ 4

$ 0

Debt

Fixed assets and net working capital

            27

$          31

            31

$          31

Equity market capitalization
(1 million shares at $31 per share)

Assume that Rational wins a lawsuit and is paid $3 million in cash. Its market capitalization rises by that amount. It decides to pay out $7 per share instead of $1 per share.

a. What happens to Rational’s stock price if the payout comes as a cash dividend?

New stock price


b. What happens to Rational’s stock price if the payout comes as a share repurchase? (Round your intermediate calculations to the nearest whole number.)

New stock price
Answer #1

a)

law suit worth = $3 million

existing Equity = $31 million

New market capitalization = 31 + 3 = $34 million

Amount paid as dividend = $7

new Price = 34 - 7 = $27

b)

share price after law suit = 31 + 3 = 34

number of shares to be bought back = 7 million / 34 = 205882.35

remaining shares = 1000000 - 205882.35 = 794117.65

value after buy back = 34 - 7 = $27 million

Share price = 27,000,000 / 794117.65 = $34

(please give thumbs up if it's helpful. thank you)